OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify sqrt(320)and provide the simplified form. medals for alll

OpenStudy (anonymous):

Find the biggest perfect square (4,16,25,36 ect ect..) that goes into 320 evenly. In this case the biggest perfect square is 64. So you can rewrite it as \[\sqrt{64}*\sqrt{5}\] and then simplify leaving you with \[8\sqrt{5}\]

OpenStudy (anonymous):

2nd method:- 320 = 32 x 10 = 64 x 5 so answer is as before (note that 10 = 5 x 2).